OpenStudy (anonymous):

simplify : 3x − 6y + 2z − 5x − 3y + 8z

OpenStudy (anonymous):

Combine like terms

OpenStudy (anonymous):

i don't see a equal sign

OpenStudy (anonymous):

3x-5x = -2x -6y-3y = -9y 2z+8z = 10z -2x-9y+10z

OpenStudy (anonymous):

there is no equal sign

OpenStudy (anonymous):

you did it correctly

OpenStudy (anonymous):

Equal sign is only there when you need it... Most of the time, the equation is set to 0, so the equal sign does not need to be there.

OpenStudy (anonymous):

so which one is the answer

OpenStudy (anonymous):

-2x-9y+10z

OpenStudy (anonymous):

thanks